# The Host Galaxies of Radio AGN: New Views from Combining LoTSS and MaNGA Observations

### v1

@@ 3.2 Radio-excess AGN and their jet luminosities
We classify a MaNGA galaxy as a radio-excess AGN host if its is at least higher than the predicted by its SFR and Equation 1. Considering the uncertainty of these measurements, the classification can be expressed as follows: . This method results in a new sample of 616 (11%) radio AGN hosts among 5578 LoTSS-MaNGA galaxies. Other detections and non-detections are considered as galaxies without secure evidence of radio excess (hereafter referred to as ‘normal galaxies’ for convenience).”
We also calculate the radio jet luminosity by subtracting the star formation contribution from the total radio luminosity: , where is the 144MHz luminosity predicted by Equation 1. The error on the , estimate is derived from the observed uncertainty in and the uncertainty in , and thus includes both and . It is important to note that some star-forming galaxies (SFGs) may have negative values if they lie below the radio-SFR relation. This is not physical and to address this, we adopt as the upper limit of for these systems. We note that the high SFR SFGs will then have higher upper limits. For non-detections, the upper limit of is set equal to the upper limit of . The consequences of these selection effects are discussed in Section 5.1.
@@
Fig. 3 illustrates our radio AGN classification for the entire LoTSS-MaNGA sample. Detections above the limit are classified as radio-excess AGN (red), while other detections (blue) and non-detections (gray) are classified as normal galaxies with an upper limit for jet luminosity. We also show the radio AGN classified in Best & Heckman (2012) using NVSS and FIRST data, i.e., the Comerford et al. (2020) sample (green stars), demonstrating not only that our sample probes radio AGN to significantly lower luminosities, but also that some high SFR SFGs may be misclassified as AGN using their methods. 77% of radio AGN in our sample are newly classified.
@@ 3.3 Control sample selection
It is well-established that galaxy age and star formation rate (SFR) are strongly correlated with stellar mass (e.g., Gallazzi et al. 2005; Guo et al. 2015). To evaluate possible different star formation histories of radio AGN host galaxies, it is essential to construct mass-controlled samples for comparison. For each radio AGN host, we create a control sample consisting of normal galaxies (no confirmed radio excess) that meet the following criteria:
Jet luminosity: upper-limit
Stellar mass: —log()— 0.3
Physical effective radius: —log()— 0.2
Angular diameter distance: —log()— 0.2
These criteria are set to select as many control sample as possible while considering the uncertainty level of the measurements.
Since radio AGN tend to be more massive than normal galaxies, there are typically more less-massive galaxies available in the control sample for a given AGN. To maintain a roughly equal median stellar mass between the AGN and control samples, for each AGN we select an equal number of more massive galaxies and less massive galaxies that meet the above criteria. Out of the 616 radio AGN, only three lack any control galaxies that meet these criteria and are discarded.
@@ 4.1 Radio AGN prefer massive quiescent hosts
Fig. 5 illustrates the global star formation rate of our RDAGN sample. The majority of RDAGN hosts (represented by red stars) are situated below the star-forming galaxy main sequence (SFMS, depicted by the black solid line). Galaxies below the 3 lower limit of the SFMS are defined as ‘quenched’, as they have transitioned from the SFMS to the quiescent population. Notably, the RDAGN hosts are predominantly composed of quenched galaxies (567 of 616, 92%). Furthermore, we investigated the occurrence of RDAGN in the quenched galaxy population. As shown in the upper panel of Fig. 5, the fraction of RDAGN among all quenched galaxies, , exhibits an increasing trend with stellar mass, reaching at least 50% at . We note here that this fraction can only be considered as a lower limit at the LoTSS sensitivity, because LoTSS is still not deep enough to confirm if some weak radio excess exists in these non-detected galaxies. Deeper radio continuum surveys in the future will probe fainter radio AGN.
@@
These results are consistent with our current knowledge about RDAGN in the local universe. Previous studies have indicated a preference for RDAGN to be hosted in massive, old, and elliptical galaxies (Best et al. 2005a). Recent LOFAR data also suggest that all the massive galaxies have a RDAGN inside (Sabater et al. 2019). The fraction observed in our sample is considerably smaller (41%), but is still reasonable considering that we use very different RDAGN classifications and the fraction definition compared to theirs. These results all suggested that the radio AGN hosts currently have old stellar populations and have a early and fast assembly history, which we will investigate in detail in Section 4.2.
@@
In Fig. 6, we explore if there is a relationship between global quenching and the power of radio jets. We quantify the degree of quenching by measuring how far galaxies lie below the SFMS: , i.e. smaller (more negative) indicates a more quenched galaxy. Galaxies below the 3 threshold of the SFMS (the green dashed line) are classified as quenched. The upper panel of Fig. 6 displays the fraction of quenched galaxies among all RDAGN hosts () across different bins. Due to the limits of measurements, our sample is complete only when the RDAGN are more luminous than . At this jet luminosity, 95% (140 out of 147) of RDAGN host galaxies are quenched, indicating that nearly all hosts of luminous RDAGN are quenched. For less luminous RDAGN, the range of (the shadow region in the upper panel) is determined by considering two extreme assumptions. First, as the lower limit that all upper limits indicate that their hosts do not host RDAGN. Second, as the upper limit that all hosts host radio jets with below the upper limits. Even under the first assumption, the fraction remains at least 50% down to AGN luminosities of . These fractions support the hypothesis that there are connections between the existence of RDAGN and the quiescence of their hosts. However, the Pearson coefficient between and is approximately -0.1, indicating that the quenching level is irrelevant to the radio AGN power.
@@
We note that there are some galaxies with upper limits in Fig. 6 (blue arrows). This is because star formation dominates the total radio luminosity in those galaxies, making it hard to classify weak radio AGN based on radio luminosity. For these galaxies, the LoTSS sensitivity is no longer the limitation to confirm if there is radio AGN activity. Higher resolution star formation and radio continuum maps would be needed to decompose the AGN jets from star formation at sub-galactic scales.
In summary, quiescent galaxies dominate the RDAGN hosts, with increasing with while shows no correlation with . These results suggest that radio AGN feedback is more likely to maintain galaxy quiescence rather than drive the galaxy quenching from the SFMS, which is consistent with Mulcahey et al. (2022).
@@
‘Radio-mode’ AGN feedback is believed to be necessary to regulate the cooling flows in dense cluster environment and help massive galaxies maintain quiescent (e.g., Fabian 2012). Radio-mode AGN feedback models have been implemented in cosmological hydrodynamical simulations, where the existence of jet mode feedback is crucial for preventing galaxies from becoming too massive (Torrey et al. 2014). Because the jets dump most of their energy at large distances from the galaxy and because the radio jet power varies on very short timescales compared to the cooling time of the gas in dark matter halos, the luminosity of the jet may not correlate very strongly with the star formation rate in the galaxy. In addition, several observational studies have found weak correlations between the radio jet luminosity and both the kinetic power of the jet and the SMBH accretion rate (e.g., Fabian 2012; Russell et al. 2013), suggesting that the energy transfer from radio jets to surrounding gas is inefficient. These factors may explain why the radio luminosity of our RDAGN sample is irrelevant to the quenching level of the host galaxy.
@@ 4.2 Radio AGN host galaxies quenched earlier
Fig. 7 shows the comparison of median mass assembly histories between RDAGN hosts and their controls. For both radio AGN hosts (red solid lines) and their controls (blue dashed lines), galaxies form the central region (0-0.5 ) earlier than the middle region (0.5-1 ) and outskirts (1-1.5 ), following an ‘inside-out’ formation scenario. At all radii, RDAGN built their mass faster and quenched earlier, with the early quenching more pronounced in the central region. Note that this difference does not necessarily imply a direct connection between RDAGN and the evolution of their hosts, because of the very different timescales of radio jet evolution compared to the buildup of the stellar mass. It is possible that similar conditions prevail during the bulk of the stellar mass build-up of both the radio AGN host galaxies and the control sample galaxies. Faster growth could be a consequence of radio AGN being located in denser environments (e.g., Peng et al. 2010). This scenario is also favored by the more significant impact in the central region; in the standard CDM cosmological paradigm where structure forms hierarchically, the formation time of the inner region of the galaxy is boosted to earlier epochs in dense environments. RDAGN activity may only be triggered after their hosts exceed a certain mass threshold and may only play an important role in maintaining the quiescence at a late stage of the host galaxies’ evolution. Such a scenario is in keeping with the implementation of radio mode feedback in semi-analytical models of galaxy formation (e.g., Croton et al. 2006).
@@
4.3 The impact of radio AGN on ISM revealed by the spatially resolved spectra
@@
To investigate which population of RDAGN are contributing to the stacked emission line excess and deficiency, we divide the RDAGN sample into different subsamples by jet luminosity, radio morphology, and stellar mass, respectively. In each category, we create 16 stacked different spectra from these four subsamples and from four radial bins (0-0.5 , 0.5-1 , 1-1.5 , ). We then focus on the wavelength region between 6380 and 6670 to quantify the width and strength of the H+[N II] features. We subtract the continuum by fitting it polynomially and then use the single component Gaussian model to fit the equivalent width and dispersion of these three lines. Fig. 10 illustrates this approach when we stack the spectra from RDAGN with different . RDAGN are divided into four bins: , -, -, and . There are 173, 165, 130, and 144 RDAGN in these four luminosity bins, respectively. Panels from top to bottom represent sub-samples of four bins, from powerful RDAGN to weak RDAGN. Within each luminosity bin, the differential spectra are plotted for four radial bins (0-0.5 , 0.5-1 , 1-1.5 , ) from left to right. We note that the outermost region, , is based on incomplete samples because of the field of view limit of MaNGA IFUs. The equivalent width and the signal-to-noise ratio of H are listed in each panel, where positive and negative values represent stronger or weaker H in RDAGN hosts, respectively. It can be seen that stacking can reveal equivalent width differences of less than 0.5.
@@
We use a similar approach for the other two categories, radio morphology and stellar mass, and summarize the equivalent width measurements in Fig 11. The three panels from top to bottom show the results when dividing the RDAGN host sample by , radio morphology, and stellar mass. The calculation of equivalent width () is based on the same method as in DAP (Belfiore et al. 2019). For radio morphology, we visually classify the RDAGN into FR I, FR II, diffuse, and unresolved morphology based on the 6 resolution LoTSS images. The number of RDAGN in each morphology category is 78, 93, 43, and 404, respectively. In the lowest panel, we divide the RDAGN hosts into four stellar mass bins (, , , and ), which include 132, 145, 186, and 150 galaxies, respectively. Fig 11 shows the radial profiles of the line excess. Among different subsamples, the lower mass RDAGN (median : ) and compact radio sources () contribute most of the H+[N II] excess in the nuclear regions. This indicates that the the connection between radio AGN activities and optical emission is more common when the RDAGN hosts are less massive, and only exists in small regions in the galaxy center. In the outskirts (), we find emission line deficiency, mainly contributed by the low-power RDAGN population, which is possibly related to different star formation conditions.
@@
In Fig 12, we plot the distribution of the nuclear [N II]/H versus of all MaNGA galaxies using a logarithm scale (background gray contours). Emission lines from H II regions and AGN or shocks are marked by blue and red contours, which are used to classify galaxies according to location in the BPT diagram (e.g., Kauffmann et al. 2003). The red arrow indicates the direction where AGN or shock ionization becomes more dominant. We measure the the line ratio and of the emission excess for the whole RDAGN sample in the nuclear bin and the deficiency in the outskirts, and plot them as a green circle and square respectively. We note that the measurements are shown as upper limits, because the stacking procedure will bring an unknown level of line broadening. The central line excess shows a high [N II]/H ratio and is located in the region predicated by AGN or shock models (e.g., D’Agostino et al. 2019). This supports the hypothesis that the central RDAGN is heating and ionizing the nuclear gas. However, the excess is weak () and not correlated with jet luminosity, e.g., the most powerful RDAGN population () do not show significant line excess. This suggests that the RDAGN have a weak and complex impact on the ISM properties in the central region of the galaxy. It is worth noting that the resolution of the LoTSS images and MaNGA IFUs is at the kiloparsec level, necessitating future high-resolution observations to confirm whether stronger effects are seen at smaller distances from the SMBH. In addition, high-resolution images are also key to distinguishing whether the emission is from the jet-related shock or from AGN photoionization.
Compared to the central region, the deficiency in the outskirts is located much closer to the star-forming region. This deficiency seems to be stronger in the lower luminosity RDAGN sub-samples, so weaker star formation in the outskirts of RDAGN hosts is a likely explanation. We note that the lower luminosity RDAGN hosts are more likely to be found at lower redshifts due to the survey sensitivity, and the fact that they are found in less massive host galaxies is partially a selection effect. Nevertheless, the significant absorption-like features in the low power RDAGN host differential spectra still suggest that in galaxies, RDAGN hosts have weaker outer region star formation than their controls.

### v3

@@ 3.2 Radio-excess AGN and their jet luminosities
We classify a MaNGA galaxy as a radio-excess AGN host if its is at least level higher than the predicted by its SFR and Equation 1. Considering the uncertainty of these measurements, the classification can be expressed as follows: . This method results in a new sample of 616 (11%) radio AGN hosts among 5548 LoTSS-MaNGA galaxies. This way the AGN sample should all be secure AGN, but it may be incomplete at lower radio luminosities. Other detections and non-detections are considered as galaxies without secure evidence of radio excess (hereafter referred to as ‘non-RDAGN’ for convenience).”
We also calculate the radio jet luminosity by subtracting the star formation contribution from the total radio luminosity: , where is the 144MHz luminosity predicted by Equation 1. The error on the , estimate is derived from the observed uncertainty in and the uncertainty in , and thus includes both and .
For non-RDAGNs, we need an estimate of the upper limit of to build a reliable control sample. The SMBHs in non-RDAGN may still have weak radio activities hidden in the host star formation555For example, we can consider a galaxy which has with a radio AGN inside it. It will be classified as non-RDAGN because the radio AGN is hidden in the star formation. This way it should not be selected as a control galaxy of a RDAGN with , and the upper limit can help us to quantify this effect.. For non-RDAGN with radio detections, we consider that the radio continuum is dominated by star formation rather than AGN, thus we assume a maximum upper limit of . For non-RDAGN not detected by LoTSS, we adopt the upper limit of radio luminosity (defined in Section 2.3) also as the upper limit of . By these definitions, the non-RDAGN’s upper limit is the maximum possible luminosity of the ‘hidden AGN’. A non-RDAGN with the upper limit much lower than the of a RDAGN can thus be confidently selected as a control galaxy. We note that the high SFR star-forming galaxies (SFGs) with radio detection will then have higher upper limits, because they can have more radio luminosity dominated by star formation than some low SFR RDAGN. The consequences of this selection effect are discussed in Section 5.1.
@@
Fig. 3 illustrates our radio AGN classification for the entire LoTSS-MaNGA sample. Detections above the limit are classified as radio-excess AGN (red), while other detections (blue) and non-detections (green) are classified as non-RDAGN with an upper limit for jet luminosity. We also show the radio AGN classified in Best & Heckman (2012) using NVSS and FIRST data, i.e., the Comerford et al. (2020) sample (black stars), demonstrating not only that our sample probes radio AGN to significantly lower luminosities, but also that some high SFR SFGs may be misclassified as AGN using their methods. 77% of radio AGN in our sample are newly classified.
, respectively. Detections with observed larger than the limit are classified as radio-excess AGN (red stars). Other detections (blue) and non-detections (green) are considered as radio normal galaxies (non-RDAGN). We plot the matched radio AGN classified in Best & Heckman (2012) and Comerford et al. (2020) as a comparison (black stars). Our sample includes more low luminosity radio AGN and can distinguish some high SFR SFGs, and 77% of radio AGN in our sample are newly classified.
@@ 3.3 Control sample selection
It is well-established that galaxy age and star formation rate (SFR) are strongly correlated with stellar mass (e.g., Gallazzi et al. 2005; Guo et al. 2015). To evaluate possible different star formation histories of radio AGN host galaxies, it is essential to construct mass-controlled samples for comparison. For each radio AGN host, we create a control sample consisting of non-RDAGN (no confirmed radio excess) that meet the following criteria:
Jet luminosity: upper-limit,
Stellar mass: —log()— 0.3,
Physical effective radius: —log()— 0.2,
Angular diameter distance: —log()— 0.2.
The first criterion ensures that even if the control galaxy contains an unclassified radio AGN , its maximum jet luminosity is still marginal compared to that of the corresponding RDAGN (at most 20%). Other criteria are set to select as many control sample as possible while considering the uncertainty level of the measurements. For instance, the stellar mass calculation results from different methods are found to have a scatter of 0.2-0.3 dex (e.g. Małek et al. 2018; Neumann et al. 2022; Sánchez et al. 2022), the scatters among the fitted from different profiles (e.g. Petrosian and Sérsic) are about 0.1-0.2 dex.
Since radio AGN tend to be more massive than non-RDAGN, there are typically more less-massive galaxies available in the control sample for a given AGN. To maintain a roughly equal median stellar mass between the AGN and control samples, for each AGN we select an equal number of more massive galaxies and less massive galaxies that meet the above criteria. Out of the 616 radio AGN, only three lack any control galaxies that meet these criteria and are discarded.
@@ 4.1 Radio AGN prefer massive quiescent hosts
Fig. 5 illustrates the global star formation rate of our RDAGN sample. The majority of RDAGN hosts (represented by red stars) are situated below the star-forming galaxy main sequence (SFMS, depicted by the black solid line). Galaxies below the 3 lower limit of the SFMS are defined as ‘quenched’, as they have transitioned from the SFMS to the quiescent population. Notably, the RDAGN hosts are predominantly composed of quenched galaxies (567 of 616, 92%). Furthermore, we investigated the occurrence of RDAGN in the quenched galaxy population. As shown in the upper panel of Fig. 5, the fraction of RDAGN among all quenched galaxies, , exhibits an increasing trend with stellar mass, reaching at least 50% at , which indicates increased recurrence of RDAGN in massive galaxies. We note here that this fraction can only be considered as a lower limit at the LoTSS sensitivity, because LoTSS is still not deep enough to confirm if some weak radio excess exists in these non-detected galaxies. Deeper radio continuum surveys in the future will probe fainter radio AGN.
@@
These results are consistent with our current knowledge about RDAGN in the local universe. Previous studies have indicated a preference for RDAGN to be hosted in massive, old, and elliptical galaxies (Best et al. 2005a). Recent LOFAR data also suggest that locally, all the massive galaxies have a RDAGN inside (Sabater et al. 2019). The fraction observed in our sample is considerably smaller (41%), but is still reasonable considering that we use very different RDAGN classifications and the fraction definition compared to theirs. These results all suggested that the radio AGN hosts currently have old stellar populations and have an early and fast assembly history, which we will investigate in detail in Section 4.2.
@@
In Fig. 6, we explore if there is a relationship between global quenching and the power of radio jets. We quantify the degree of quenching by measuring how far galaxies lie below the SFMS: , i.e. smaller (more negative) indicates a more quenched galaxy. Galaxies below the 3 threshold of the SFMS (the black dashed line) are classified as quenched. The upper panel of Fig. 6 displays the fraction of quenched galaxies among all RDAGN hosts () across different bins.
Due to the nature of the radio luminosity function of SFGs and radio AGN (e.g. Sabater et al. 2019), the completeness of the RDAGN sample will decrease at low , where faint RDAGN in luminous SFGs will become indistinguishable. At , all the radio sources are confidently classified as RDAGN (see Figure 3), and 95% (140 out of 147) of them are below the limit of the SFMS, indicating that nearly all hosts of luminous RDAGN are quenched. For less luminous RDAGN, the range of (the shadow region in the upper panel) is determined by considering two extreme assumptions. First, as the lower limit that all upper limits indicate that their hosts do not host RDAGN. Second, as the upper limit that all hosts host radio jets with below the upper limits. Even under the first assumption, the fraction remains at least 50% down to AGN luminosities of . These fractions support the hypothesis that there are connections between the existence of RDAGN and the quiescence of their hosts. However, the Pearson coefficient between and is approximately -0.1, indicating that the quenching level is irrelevant to the radio AGN power.
@@
We note that some radio-detected galaxies have upper limits in Fig. 6 (blue arrows). This is because star formation dominates the total radio luminosity in those galaxies, making it hard to classify weak radio AGN based on radio luminosity. For these galaxies, the LoTSS sensitivity is no longer the limitation to confirm if there is radio AGN activity. Higher resolution star formation and radio continuum maps would be needed to decompose the AGN jets from star formation at sub-galactic scales.
@@
In summary, quiescent galaxies dominate the RDAGN hosts, with increasing with while shows no correlation with . ‘Radio-mode’ AGN feedback is believed to be necessary to regulate the cooling flows in dense cluster environment and help massive galaxies maintain quiescent (e.g., Fabian 2012). Radio-mode AGN feedback models have been implemented in cosmological hydrodynamical simulations, where the existence of jet mode feedback is crucial for preventing galaxies from becoming too massive (Torrey et al. 2014). Because the jets dump most of their energy at large distances from the galaxy and because the radio jet power varies on very short timescales compared to the cooling time of the gas in dark matter halos, the luminosity of the jet may not correlate very strongly with the star formation rate in the galaxy. In addition, several observational studies have found weak correlations between the radio jet luminosity and both the kinetic power of the jet and the SMBH accretion rate (e.g., Fabian 2012; Russell et al. 2013), suggesting that the energy transfer from radio jets to surrounding gas is inefficient. These factors may explain why the radio luminosity of our RDAGN sample is irrelevant to the quenching level of the host galaxy.
@@ 4.2 Radio AGN host galaxies quenched earlier
Fig. 7 shows the comparison of median mass assembly histories between RDAGN hosts and their controls. For both radio AGN hosts (red solid lines) and their controls (blue dashed lines), galaxies form the central region (0-0.5 ) earlier than the middle region (0.5-1 ) and outskirts (1-1.5 ), following an ‘inside-out’ formation scenario. At all radii, RDAGN built their mass faster and quenched earlier, with the early quenching more pronounced in the central region. Note that this difference does not necessarily imply a direct connection between RDAGN and the evolution of their hosts, because of the very different timescales of radio jet evolution compared to the buildup of the stellar mass. It is possible that different conditions prevail during the bulk of the stellar mass build-up in radio AGN host galaxies and the control galaxies. Faster growth could be a consequence of radio AGN being located in denser environments (e.g., Peng et al. 2010). This scenario is also favored by the more significant impact in the central region; in the standard CDM cosmological paradigm where structure forms hierarchically, the formation time of the inner region of the galaxy is boosted to earlier epochs in dense environments. RDAGN activity may only be triggered after their hosts exceed a certain mass threshold and may only play an important role in maintaining the quiescence at a late stage of the host galaxies’ evolution. Such a scenario is in keeping with the implementation of radio mode feedback in semi-analytical models of galaxy formation (e.g., Croton et al. 2006).
@@
4.3 Emission line excess in the nuclei of radio AGN hosts
@@
To investigate which population of RDAGN are contributing to the stacked emission line excess and deficiency, we divide the RDAGN sample into different subsamples by jet luminosity, radio morphology, and stellar mass, respectively. In each category, we create 16 stacked different spectra from these different subsamples and from four radial bins (0-0.5 , 0.5-1 , 1-1.5 , ). We then focus on the wavelength region between 6380 and 6670 to quantify the width and strength of the H+[N II] features. We subtract the continuum by fitting it polynomially and then use the single component Gaussian model to fit the equivalent width and dispersion of these three lines.
In order to ensure that our spectral stacks of radio AGN are fully complete, we use the 90% completeness cut in radio luminosity as a function of redshift to divide the radio AGN sample into four luminosity bins constrained by redshift. The median AGN luminosities for the bins are chosen to be , , , and W Hz-1. In total, 388 RDAGN are left to be used for the spectral stacking analysis.
Fig. 11 illustrates this approach when we stack the spectra from RDAGN with different . We set luminosity bins to make the median values equal to , and . There are 70, 122, 121, and 75 RDAGN in these four luminosity bins, respectively. Panels from top to bottom represent subsamples of four bins, from powerful RDAGN to weak RDAGN. Within each luminosity bin, the differential spectra are plotted for four radial bins (0-0.5 , 0.5-1 , 1-1.5 , ) from left to right. We note that the outermost region, , is based on incomplete samples because of the field of view limit of MaNGA IFUs. The equivalent width and the signal-to-noise ratio of H are listed in each panel, where positive and negative values represent stronger or weaker H in RDAGN hosts, respectively. It can be seen that stacking can reveal equivalent width differences of less than 0.5.
@@
We use a similar approach for the other two categories, radio morphology and stellar mass, and summarize the equivalent width measurements in Fig 12. The three panels from top to bottom show the results when dividing the RDAGN host sample by , radio morphology, and stellar mass. The calculation of equivalent width () is based on the same method as in DAP (Belfiore et al. 2019). For radio morphology, we visually classify the RDAGN into FR I, FR II, diffuse, and compact (or unresolved) morphology based on the 6 resolution LoTSS images. The number of RDAGN in each morphology category is 81, 77, 38, and 192, respectively. In the lowest panel, we divide the RDAGN hosts into four stellar mass bins with median of , and , which include 47, 50, 145, and 146 galaxies, respectively. Fig 12 shows the radial profiles of the line excess. Among different subsamples, the compact radio sources () show strong emission line excess, consistent with the scenario that the compact radio AGN are younger (e.g. Chilufya et al. 2024) and therefore have a greater impact on gas ionization around the central SMBHs. In addition, the lower mass RDAGN (median : ) have the strongest H+[N II] excess in their nuclear regions. This indicates that the connection between radio AGN activity and optical emission is more common when the RDAGN hosts are less massive. In the outskirts (), we find weak emission line deficiency, mainly contributed by resolved RDAGN (FR I, FR II, and diffuse RDAGN). These are likely to be older sources where the jet impacts the ISM at larger distances.
of the excess or deficiency () in different RDAGN subsamples. Panels from top to bottom show the results when dividing RDAGN by jet luminosity, radio morphology, and stellar mass, respectively. Profiles of different subsamples are shown in different colors and markers. Positive represents that RDAGN have stronger emission lines and vice versa. Low luminosity (), compact (), and lower mass RDAGN () contribute most to the nuclear excess. Weak emission line deficiency is found in the outskirts of large size RDAGN population.
@@
In Fig 13, we plot the distribution of the nuclear [N II]/H versus of all MaNGA galaxies using a logarithm scale (background gray contours). Emission lines from H II regions and AGN or shocks are marked by blue and red contours, which are used to classify galaxies according to location in the BPT diagram (e.g., Kauffmann et al. 2003). The red arrow indicates the direction where AGN or shock ionization becomes more dominant. We measure the the line ratio and of the emission excess for the whole RDAGN sample in the nuclear bin and the deficiency in the outskirts, and plot them as a green circle and square respectively. We note that the measurements are shown as upper limits, because the stacking procedure will bring an unknown level of line broadening. The central line excess shows a high [N II]/H ratio and is located in the region predicated by AGN or shock models (e.g., D’Agostino et al. 2019). This supports the hypothesis that the central RDAGN is heating and ionizing the nuclear gas. However, the excess is weak () and not correlated with jet luminosity. This suggests that the RDAGN have a weak and complex impact on the ISM properties in the central region of the galaxy. It is worth noting that the resolution of the LoTSS images and MaNGA IFUs is at the kiloparsec level, necessitating future high-resolution observations to confirm whether stronger effects are seen at smaller distances from the SMBH. In addition, high-resolution images are also key to distinguishing whether the emission is from the jet-related shock or from AGN photoionization. Compared to the central region, the deficiency in the outskirts is located much closer to the star-forming region, suggesting that the RDAGN hosts tend to be more quiescent in outer regions.
In summary, we find evidence of emission line excess in the nuclei of RDAGN host galaxies. The excess is stronger in lower radio luminosity, more compact (smaller size of radio emission), and lower stellar mass RDAGN. The line excess is likely linked to the interaction of the jet with the gas in the central regions of the host galaxies of the RDAGN. The line ratios suggest AGN photoionization or shock ionization, and the excess is stronger towards the galaxy center. We note that the level of emission line excess in our sample is not correlated with jet power in the range W Hz-1. In the lowest luminosity bin, stronger excess is seen.
We note that the sample is not large enough to split the sample into a bin of more than one quantity. In order for the radio jet to produce an emission line excess, the host galaxy must contain gas. Our radio-selected sample, however, includes many massive galaxies that lack gas to interact with the radio jets. This may explain the stronger line excess observed in lower mass RDAGN, as lower mass galaxies usually contain more gas (e.g. Saintonge & Catinella 2022). The excess observed in the lowest radio luminosity bin may also result from the fact that these are mainly found in low mass galaxies.